# RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Lifelong Learning in Physical Embodied Systems

### v3

@@ Abstract
We present RoboMemory, a brain-inspired multi-memory framework for lifelong learning in physical embodied systems, addressing critical challenges in real-world environments: continuous learning, multi-module memory latency, task correlation capture, and infinite-loop mitigation in closed-loop planning. Grounded in cognitive neuroscience, it integrates four core modules: the Information Preprocessor (thalamus-like), the Lifelong Embodied Memory System (hippocampus-like), the Closed-Loop Planning Module (prefrontal lobe-like), and the Low-Level Executer (cerebellum-like) to enable long-term planning and cumulative learning. The Lifelong Embodied Memory System, central to the framework, alleviates inference speed issues in complex memory frameworks via parallelized updates/retrieval across Spatial, Temporal, Episodic, and Semantic submodules. It incorporates a dynamic Knowledge Graph (KG) and consistent architectural design to enhance memory consistency and scalability. Evaluations on EmbodiedBench show RoboMemory outperforms the open-source baseline (Qwen2.5-VL-72B-Ins) by 25% in average success rate and surpasses the closed-source State-of-the-Art (SOTA) (Claude3.5-Sonnet) by 5%, establishing new SOTA. Ablation studies validate key components (critic, spatial memory, long-term memory), while real-world deployment confirms its lifelong learning capability with significantly improved success rates across repeated tasks. RoboMemory alleviates high latency challenges with scalability, serving as a foundational reference for integrating multi-modal memory systems in physical robots.
Figure 1: The brain-inspired architecture of the RoboMemory resembles the biological nervous system. It maps biological neural components, enabling the agent to interact with diverse environments (real-world, Habitat, ALFRED) and robotic hardware for long-term planning and lifelong learning.
@@ 1 Introduction
Driven by the rapid advancements in Vision-Language Models (VLMs) hurst2024gpt ; bai2025qwen2 , VLM-based agents have been increasingly deployed in embodied tasks, leveraging pre-trained knowledge and multimodal understanding capabilities to interact with physical environments park2023generative ; hu2023toward .
Current research primarily focuses on optimizing performance in single tasks within virtual simulators or controlled real-world setups belkhale2024rt ; ahn2022can ; yang2024guiding , or relies on oversimplified memory frameworks. These frameworks lack mechanisms to model interdependencies between distinct tasks. In practice, real-world embodied agents must operate over a lifetime, handling sequential tasks where prior experiences (e.g., learning to “open a fridge” to retrieve an apple) directly shape subsequent performance (e.g., efficiently accessing other objects in similar containers). This demands long-term memory systems that not only retain experiences but also capture cross-task influences, enabling cumulative improvement rather than isolated task execution.
Prior efforts to integrate memory systems into embodied frameworks tan2024cradle ; glocker2025llm ; wang2023voyager ; agashe2024agent ; fu2024msi have advanced long-term planning and lifelong learning. However, most of them concentrate on virtual environments and fail to generalize to the real world. While the embodied frameworks that are designed for the real world yuan2025being ; shi2025hi lack critical memory modules. But on the other hand, a complex memory architecture will suffer from prohibitive latency, which makes it hard to use.
To address these gaps, we propose RoboMemory, a brain-inspired multi-memory framework explicitly designed for lifelong learning in the real world. This highly parallelized, hierarchical architecture enables long-term planning and lifelong learning in real-world environments. Drawing on cognitive neuroscience milner1998cognitive , it comprises four core components (Fig. 1): an Information Preprocessor (thalamus-like) for multimodal integration; a Lifelong Embodied Memory System (hippocampus-like), the core part that adopts a three-tier structure (long-term, short-term, working memory) grounded in cognitive neuroscience to organize experiential data and object spatial relationships, with a parallelizable memory paradigm unifying information updating and retrieval across modules to mitigate latency; a Closed-Loop Planning Module (prefrontal lobe-like) for high-level action sequencing; and a Low-level Executer (cerebellum-like). For robust real-world deployment, RoboMemory employs a dual-system architecture yuan2025being ; shi2025hi where the upper layer (embodied agent) outputs abstract high-level actions, and a Vision-Language-Action (VLA) model with Simultaneous Localization and Mapping (SLAM) system translates these into low-level commands executable by robots.
We validated RoboMemory (excluding the executor) in EmbodiedBench, a long-horizon planning benchmark environment yang2025embodiedbench . Results demonstrate substantial improvements: with Qwen2.5-VL-72b as the backbone, RoboMemory increases average success rates by 25% over the base model and 5% over the closed-source state-of-the-art (SOTA) Claude-3.5-sonnet. Moreover, we test RoboMemory in a real-world environment, where RoboMemory executed 15 diverse tasks twice (one for learning and one for testing). By validating the RoboMemory’s performance between two executions, we verify that the RoboMemory has basic lifelong learning ability in the real world. In addition, we do ablation studies to quantify component-level contributions, and a comprehensive error analysis contextualizes limitations. In conclusion, our contributions fall into three aspects:
  • Inspired by the brain’s unified memory mechanisms, we design a lifelong embodied memory system with four parallel modules (Spatial, Temporal, Episodic, and Semantic) under a unified framework. This framework supports parallelized update and retrieval across modules, mitigating latency accumulation in complex systems while facilitating coherent knowledge integration for lifelong learning.
  • We propose a retrieval-based incremental Knowledge Graph (KG) update algorithm for dynamic spatial memory evolution. It enables efficient, consistent updates by retrieving relevant subgraphs, detecting local conflicts, and merging new information, alleviating scalability bottlenecks of traditional incremental strategies in dynamic environments.
@@
  • RoboMemory enables lifelong learning in real-world physical robots: it executes sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements. This demonstrates practical long-term autonomous learning in physical scenarios, with reduced reliance on simulated pre-training.
2 Related Work
2.1 VLM/LLM-based Agentic Frameworks in Embodied Tasks
The rapid advancement of VLMs/LLMs has led to diverse agent frameworks in embodied environments yao2022react ; song2023llm ; lin2024swiftsage . Embodied tasks involve partial observability and long-horizon planning, requiring memory systems to retain context. Some use time-ordered context buffers for short-term memory (due to VLMs/LLMs’ limited long-context processing) yao2022react ; packer2023memgpt ; others adopt experience buffers as long-term semantic memory fu2024msi ; shinn2024reflexion . For long-duration tasks, skill libraries serve as procedural memory, with agents accumulating skills via interaction wang2023voyager ; tan2024cradle . Recent efforts integrate diverse memories zhang2023building ; tan2024cradle ; agashe2024agent but focus on virtual/GUI environments, leaving real-world multi-modal memory support for long-term planning under-explored.
2.2 Vision Language Action Model
Current work on VLA models uses imitation learning to output low-level controls from language and visuals black2024pi_0 ; zhao2023learning ; bjorck2025gr00t ; kim2024openvla but is limited to tabletop tasks and single actions, restricting long-horizon planning. VLAs lack long-term execution abilities, while high-level agents excel at planning. Recent works combine high-level frameworks with VLA executors, some augmented with simple memory shi2025hi ; tan2025roboos ; yuan2025being ; yang2025agentic for longer tasks. However, real-world robots need more sophisticated memory to handle continuous multi-task operations over extended periods.
2.3 Memory Frameworks
Many previous works improve long-term planning via memory systems: Voyager wang2023voyager uses a skill library in Minecraft but lacks diverse memory types; CoELA zhang2023building includes procedural, semantic, and episodic memory with a task-specific 2D map; MSI-Agent fu2024msi utilizes insight as long-term memory for in-task learning. Hippo Retrieval Augmented Generation (RAG) gutierrez2024hipporag mimics the hippocampus and introduces KGs as long-term memory indices burgess2002human ; chen2020review , enhancing retrieval. However, the previous approach is mainly focused on constructing a KG with static long context, such as a book, but it is hard to update the graph. We need to update the information in KG for the embodied task. Our approach builds a more general LLM-based memory system using a dynamic KG like Hippo RAG, which is designed for embodied tasks.
Figure 2: RoboMemory architecture with working pipeline and memory mechanisms. (a) Left: The agent’s pipeline. Parallel Step Summarizer and Query Generator in Information Processor (1) generate updates/queries for Lifelong Embodied Memory (2). These memories enable Closed-Loop Planning (3) for tasks like “slice and pick up the apple”—the Planner generates plans, while the Critic and memories adjust decisions via feedback from visual inputs/results (4). (b) Right: Spatial and Semantic memories operate in parallel with isomorphic updates. Internally, Spatial memory maintains a relevance/similarity-updated KG, and Semantic memory manages a Vector DB with analogous logic.
@@ 3 RoboMemory
As illustrated in Fig. 2, the key design of RoboMemory is a unified memory paradigm. This paradigm aims to streamline memory operations: during updates, only memory items relevant to new information are targeted, while retrieval leverages rule-based information gathering to enhance efficiency. In the following sections, we detail the specific design of each component and how they integrate within this framework to enable real-world lifelong learning.
@@ 3.1 Information Preprocessor
For each time-step $i$, RoboMemory receives a visual observation $\mathcal{O}_{i}$—a single RGB frame in simulation or a short video snippet from the onboard camera on the real robot—capturing what the agent sees while acting. An Information Preprocessor constitutes the system’s perceptual front-end, converting this multimodal input into text that can be indexed and searched. To keep latency low, the preprocessor runs two lightweight modules in parallel:
  1. 1.
    Step summarizer $\mathcal{S}$: Transforms $\mathcal{O}_{i}$ into a concise textual description $s_{i}$ of the just-executed action. The string $s_{i}$ is stored as the system’s working memory.
  2. 2.
    Query generator $\mathcal{Q}$: Derives a query $q_{i}$ from the same observation $\mathcal{O}_{i}$, which is used to probe long-term memory for relevant episodes.
Together, $\mathcal{S}$ and $\mathcal{Q}$ provide a swift, text-based interface between raw sensory data and RoboMemory’s retrieval machinery.
3.2 Lifelong Embodied Memory System
RoboMemory incorporates a lifelong memory system with four modules: Spatial, Temporal, Episodic, and Semantic. The four-module structure is designed to enable continuous learning in dynamic real-world environments. To mitigate latency from repeated VLM invocations, we unify the four modules’ update and retrieval processes into a single paradigm with parallel implementation. As shown in Fig. 2, all four modules perform updates and retrievals in parallel. Thus, the framework avoids latency escalation despite multiple memory components.
Memory modules in Lifelong Embodied Memory System have three types of update frequency:
  1. 1.
    Action-level, which is updated once per action;
  2. 2.
    Task-level, which is updated only after the completion of each task;
  3. 3.
    Mixed-level, which is updated at both the action level and the task level.
0: New triplet set $T_{new}$, main knowledge graph $G=(V,E)$, latest retrieval entities $Entity_{l}$, entity embeddings $\mathcal{E}:V\rightarrow\mathbb{R}^{d}$, similarity threshold $\theta$, conflict detection $VLM_{dt}$, retrieval function $\text{Retrieve}(\cdot)$
0: Updated consistent knowledge graph $G^{\prime}$
1: Step 1: Retrieve relevant subgraph based on new triplets
[…]
3: Step 2: Create local working subgraph
[…]
5: Step 3: Conflict detection within local subgraph
[…]
7: Step 4: Generate updated subgraph
[…]
9: Step 5: Merge updated subgraph back to main graph
[…]
11: Update vector database $V$ with new embeddings
12:return  $G^{\prime}$
@@ 3.2.1 Spatial-Temporal Memory System
To adapt to dynamic real-world environments, we design the Spatial-Temporal Memory. The spatial and temporal memory update in action-level frequency. The Temporal memory is a first-in-first-out (FIFO) buffer whose size is $N$. The buffer can store $N$ summary steps $s_{[i:i+N]}$. When the buffer is full, it will be fully cleared. Then, we use an LLM to summarize the short-term memories in the discarded $N$ step into a single entity, inserted as the first entry of the buffer.
The Spatial Memory is a dynamically updated KG-based module that addresses LLMs’ limitations in implicitly extracting spatial information from temporal memory. This system dynamically records spatial relationships in the scene while efficiently maintaining a dynamic KG. Unlike previous KG construction algorithms gutierrez2024hipporag ; edge2024local , which are designed for static information gathering. Our spatial memory needs to be updated frequently. To speed up the KG update process, we design a two-phase progressive approach: (1) Rapid Response Phase: New information is quickly buffered to avoid losing observations. (2) Local Integration Phase: On buffer saturation or conflict detection, affected local subgraphs are integrated (entity de-duplication, relationship merging, conflict detection).
This algorithm limits KG updates to the environment. For each update, we only consider relevant segments, controlling efficiency and updater context length for better performance. The algorithm is shown in Algorithm 1. Meanwhile, to illustrate the growth of dynamic spatial memory, we provide a sample in Appendix D.
Our retrieval-based incremental KG update algorithm has provable efficiency guarantees. Specifically, for a KG with $n$ nodes and maximum degree $D$, the number of nodes processed in each update is bounded by $O(D^{K})$ where $K$ is the retrieval hop distance (see Appendix C for formal proof). This ensures scalability even as the spatial memory grows over time.
3.2.2 Lifelong Learning System

### v4

# RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Interactive Environmental Learning in Physical Embodied Systems

@@ Abstract
Embodied agents face persistent challenges in real-world environments, including partial observability, limited spatial reasoning, and high-latency multi-memory integration. We present RoboMemory, a brain-inspired framework that unifies Spatial, Temporal, Episodic, and Semantic memory under a parallelized architecture for efficient long-horizon planning and interactive environmental learning. A dynamic spatial knowledge graph (KG) ensures scalable and consistent memory updates, while a closed-loop planner with a critic module supports adaptive decision-making in dynamic settings. Experiments on EmbodiedBench show that RoboMemory, built on Qwen2.5-VL-72B-Ins, improves average success rates by 25% over its baseline and exceeds the closed-source state-of-the-art (SOTA) Gemini-1.5-Pro by 3%. Real-world trials further confirm its capacity for cumulative learning, with performance improving across repeated tasks. These results highlight RoboMemory as a scalable foundation for memory-augmented embodied intelligence, bridging the gap between cognitive neuroscience and robotic autonomy.
Figure 1: RoboMemory adopts a brain-inspired architecture that maps neural components to agent modules, enabling long-term planning and interactive learning across diverse environments (real-world, Habitat, ALFRED) and robotic hardware.
@@ 1 Introduction
Recent advances in Vision-Language Models (VLMs) (hurst2024gpt, ; bai2025qwen2, ) have enabled their growing use in embodied tasks (park2023generative, ; hu2023toward, ). VLM-based embodied agents can process multimodal inputs and generate high-level textual commands (e.g., “Pick up the cup”), which require translation via tool APIs to become executable robot actions. In contrast, Vision-Language-Action models (VLAs) (kim2024openvla, ; black2024pi_0, ; bjorck2025gr00t, ; chi2023diffusion, ) produce low-level control signals directly but generally rely only on the latest observation. This limits their ability to perform long-horizon, multi-step tasks that require reasoning over task history. In summary, VLA models enable direct robot control but lack high-level planning capabilities, and VLM-based embodied agents support strategic planning but struggle with direct motor control. This highlights a key gap inherent in two distinct technical approaches to embodied intelligence.
To bridge this gap, recent work (yuan2025being, ; shi2025hi, ; tan2025roboos, ) proposes a “VLM planner + VLA executor” paradigm. Here, VLM-based embodied agents serve as high-level planners that decompose complex tasks (e.g., “make a coffee”) into executable sub-instructions (e.g., “grasp the cup”) that VLAs can complete. Although this paradigm improves performance on multi-step tasks, prior work suffers from two key limitations in real-world settings. First, real-world tasks (e.g., kitchen operations) require navigating across multiple locations to gather objects and tools, but the environment remains only partially observable at any time due to robots’ limited field of view and dynamic occlusions. This necessitates a planner with robust spatial awareness and long-term memory to maintain a consistent spatial awareness across viewpoints. However, most VLM-based agents rely on chat-style context windows (e.g., logging instruction – feedback pairs (yao2022react, )), which lack mechanisms for maintaining an overview of the environment’s spatial layout. Consequently, agents cannot reliably track object locations or recognize previously visited states. Second, pretrained VLMs are rarely trained on embodied planning trajectories, especially long-horizon, spatially grounded ones. So VLM-based agents often struggle to generalize to real-world settings (yang2025embodiedbench, ). To overcome these challenges, VLM-based planners must support interactive environmental learning — the ability to acquire, integrate, and retrieve spatial, episodic, and semantic knowledge during task execution, thereby enabling adaptation through experience.
To support long-horizon planning and interactive learning in real-world settings, agents require a comprehensive memory system with multiple specialized modules. Recent frameworks (tan2024cradle, ; glocker2025llm, ; wang2023voyager, ; agashe2024agent, ; fu2024msi, ; zhao2024expel, ; chen2024automanual, ) have integrated Retrieval-Augmented Generation (RAG)-based memory to enhance planning and interactive environmental learning (gao2023retrieval, ), but most are designed for simulated environments. A key limitation for real-world deployment is the absence of spatial memory, which is critical for building spatial awareness and providing context for planning. Additionally, existing multi-module memory systems often incur significant inference latency.
To overcome these limitations, especially the need for memory that is efficient, spatially grounded, and persistent in dynamic environments, we return to the essence of intelligence — how does the human brain plan, remember, and learn in dynamic environments? Inspired by cognitive neuroscience, we have designed RoboMemory, a parallel multi-memory architecture that simulates key functional regions of the brain. RoboMemory features a hierarchical and parallelized architecture enabling long-term planning and continuous adaptation. Drawing inspiration from cognitive neuroscience (milner1998cognitive, ), RoboMemory comprises four core components (Figure 1): (1) Information Preprocessor (thalamus-inspired) for multimodal sensory integration. (2) Comprehensive Embodied Memory System (hippocampus-inspired), which organizes experiential and spatial knowledge through a three-tier structure (long-term, short-term, and sensory memory). Within this tiered system, four memory modules: Spatial, Temporal, Episodic, and Semantic operate under a unified, parallel-update paradigm to enable coherent knowledge integration while minimizing latency. (3) Closed-Loop Planning Module (prefrontal cortex-inspired) for high-level action sequencing. These three modules provide a high-level planner with comprehensive sensory and memorization ability. (4) Low-level Executor (cerebellum-inspired), consisting of a VLA-based operation model and a SLAM-based navigation model. The Low-level Executor directly controls the robot with low-level control signals to navigate and operate in the real-world environment.
To verify whether RoboMemory truly addresses the problems of long-horizon planning and interactive learning, we evaluate RoboMemory on EmbodiedBench, a long-horizon planning benchmark (yang2025embodiedbench, ). Using Qwen2.5-VL-72B as the base model, RoboMemory improves average success rates by 25% over its base model and 3% over the closed-source state-of-the-art model, Gemini-1.5-Pro. In real-world trials, RoboMemory executed diverse tasks twice consecutively: once for environmental familiarization (learning phase) and once for memory-augmented execution (testing phase) without resetting memory. The observed performance improvement validates RoboMemory’s capacity for interactive environmental learning. We further conduct ablation studies and error analysis to quantify component contributions and identify remaining limitations. We summarize our contribution as follows:
  • We propose a brain-inspired unified embodied memory system, integrating four concurrently updated modules (Spatial, Temporal, Episodic, Semantic) into a single framework. It enables efficient, comprehensive memory operations and coherent knowledge integration, which are critical for interactive environmental learning in real-world embodied scenarios.
  • We design a retrieval-based incremental update algorithm for real-time evolution of Spatial Knowledge Graphs (KGs). By retrieving relevant subgraphs, detecting local inconsistencies, and merging new observations, it ensures efficient, consistent KG maintenance and addresses the scalability bottleneck of previous KG-based methods in embodied settings.
@@
  • RoboMemory supports interactive environmental learning for real-world physical robots: it enables sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements, demonstrating practical long-term autonomous learning in physical scenarios.
@@ 2 RoboMemory
RoboMemory is a hierarchical embodied agent system that equips robots with three core memory capabilities: historical interaction logs, dynamically updated spatial layouts, and accumulated task knowledge. As illustrated in Figure 2, each iteration, RoboMemory follows a process of “Perception – Memory – Retrieval – Planning – Execution” process, ensuring that the agent continuously calibrates its memory and behavior in dynamic environments.
First, the information preprocessor converts multimodal sensor inputs into a textual summary of the current scene, which serves as the primary input to the Comprehensive Embodied Memory. Next, the Comprehensive Embodied Memory updates its internal representations, including action histories, object locations, and experiential knowledge. After information is updated, the memory system retrieves contextually relevant entries to inform the Closed-Loop Planning Module. Then, leveraging this contextual memory, the Closed-Loop Planning Module generates high-level, text-based action instructions. Finally, these commands are dispatched to low-level executors, who will directly control the robot and complete the instructions.
Figure 2: RoboMemory architecture. (a) Left: Parallel Step Summarizer and Query Generator generate updates/queries for Comprehensive Embodied Memory. These memories enable Closed-Loop Planning for tasks like “slice and pick up the apple”—the Planner generates plans, while the Critic and memories adjust decisions via feedback from visual inputs/results. (b) Right: Spatial memory maintains a relevance/similarity-updated KG, and Semantic/Episodic memory manages a Vector DB with analogous logic. Besides, Temporal memory is implemented as a linear FIFO buffer that stores step-wise summaries generated by the Step Summarizer.
@@ 2.1 Information Preprocessor
At each time step $i$, RoboMemory receives a visual observation $\mathcal{O}_{i}$: an RGB frame (in simulation) or a short video clip (on physical robots), representing the agent’s observations.
Since raw visual data is unsuitable for direct use in memory construct and retrieval, RoboMemory first employs an information preprocessor to convert multimodal observations into textual representations, thereby providing a semantic interface for subsequent memory and planning modules. The information preprocessor executes two Vision-Language Models (VLMs) in parallel: (1) Step summarizer $\mathcal{S}$: Transforms $\mathcal{O}_{i}$ into a concise textual description $s_{i}$ of the just-executed action. The string $s_{i}$ is stored in the system’s working memory. (2) Query generator $\mathcal{Q}$: Derives a query $q_{i}$ from the same observation $\mathcal{O}_{i}$, which is used to probe long-term memory for relevant episodes.
Together, $\mathcal{S}$ and $\mathcal{Q}$ provide a swift, text-based interface between raw sensory data and provide basic information in each iteration for RoboMemory’s Comprehensive Embodied Memory System.
2.2 Comprehensive Embodied Memory System
We present a high-level memory system designed to enhance embodied agents’ long-horizon planning and enable interactive learning in dynamic, partially observable environments. Built on RAG and KG technologies (gutierrez2024hipporag, ), the system comprises Spatial-Temporal Memory and Long-Term Memory. Spatial-Temporal Memory continuously encodes object relationships and environmental layouts, as well as detailed task execution traces. These records inform Long-Term Memory, which goes through spatial-temporal information collected in each task, capturing what strategies succeeded or failed and relevant environmental facts (Figure 2).
The Comprehensive Embodied Memory module receives summarized sensory information from the information preprocessor to update both memory components in real time. Critically, all updates and retrievals are performed in parallel, ensuring efficient operation even under multiple active memory modules.
@@ 2.2.1 Spatial-Temporal Memory System
Due to the dynamic and partially observable nature of real embodied environments, completing long-term tasks requires dynamically maintaining a spatial-temporal memory system. The system records the agent’s interaction history with the environment and maintains a continuously updated spatial model of the current surroundings. To achieve these two purposes, we design spatial memory and temporal memory. The combination of spatial and temporal memory provides the embodied agent with comprehensive contextual awareness, supporting effective planning and interactive environmental learning in real-world settings.
Temporal Memory. Interactions between the robot and the environment must be recorded. For such temporally sequential memory, a simple structure suffices: a sequential buffer with automatic summarization triggered upon reaching capacity. The buffer tracks the high-level actions from the Embodied Agent, how the low-level executor actually controls the robot, and the environmental feedback received. In a specific design, temporal memory can store up to $N$ interaction summaries, each generated by an information preprocessor. When the buffer is full, rather than discarding old entries outright, we compress the oldest $N$ steps into a single summarized entry using a VLM, which is then reinserted at the front of the buffer, ensuring continuous context retention without unbounded growth.
Spatial Memory. Current spatial memory approaches often rely on RGB-D cameras to reconstruct 3D point clouds (zhang2023building, ; chang2023goat, ). However, such representations are overly detailed for high-level planning in embodied agents — precise geometric relationships (e.g., exact distances between objects) are unnecessary. Moreover, point clouds are difficult to update efficiently in dynamic scenes.
To address these problems, we advocate for an abstracted spatial memory that captures object-level entities and their spatial relationships. Inspired by gutierrez2024hipporag , we represent this memory as a KG: objects become vertices, and spatial relations are encoded as edges. The KG focuses on high-level spatial relations (e.g., “cup on table”, “key left of drawer”). Unlike geometric representations that encode precise coordinates, which are often irrelevant for high-level planning, our KG focuses on semantically meaningful, task-relevant relations. This abstraction enhances the agent’s spatial reasoning capability in dynamic environments. In our memory system, spatial memory is a dynamic KG storing high-level spatial knowledge. To overcome the limited spatial reasoning of embodied agents in dynamic environments, we use a KG.
We introduce a retrieval-driven, incremental KG update algorithm that maintains a locally modifiable, globally consistent, and dynamically adaptive spatial memory. As illustrated in the right panel of Figure 2, the update process proceeds in four steps: (1) retrieves the most relevant sub-KG around new observations. (2) Injects new relations from the current observation by a VLM-based Relation Retriever. (3) Detects and resolves conflicts between newly extracted relations and existing ones (e.g., “cup on table” vs. ‘cup in drawer”) using a VLM-based resolver, which decides whether to add, delete, or modify edges. (4) Merges back and prunes isolated vertices. Moreover, our retrieval-based incremental update algorithm comes with provable efficiency guarantees: for a KG with $n$ vertices and maximum degree $D$, the number of vertices processed per update is bounded by $O(D^{K})$, where $K$ is the retrieval hop distance (see Appendix C.3 for formal analysis). Further architectural and implementation details are provided in Appendix C.1.
2.2.2 Interactive Environmental Learning System